OpenStudy (anonymous):

What is the interquartile range of the data? 120, 140, 150, 195, 203, 226, 245, 280 A. 54 B. 90.5 C. 160 D. 235.5

OpenStudy (anonymous):

@Shay17

OpenStudy (amistre64):

q3 - q1

OpenStudy (anonymous):

so it wpold be b

OpenStudy (yanasidlinskiy):

Take a look at this: http://www.mathsisfun.com/data/quartiles.html

OpenStudy (amistre64):

since there are 8 data elements, and a quartile indicates 1/4 q = 8/4 = 2 ... which is going to mean we have to average some positional values q3 = 2(3) = 6 ... the average of the 6th and 7th positions q1 = 2(1) = 2 ... the average of the 2nd and 3rd positions

OpenStudy (amistre64):

120, 140, 150, 195, 203, 226, 245, 280 | | 130 238 so 238 - 130

OpenStudy (amistre64):

lol, 130 is not the average of 140 and 150

OpenStudy (amistre64):

238 - 145 ... assuming a got q3 right

OpenStudy (anonymous):

that's still not a choice

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=iqr%7B120%2C+140%2C+150%2C+195%2C+203%2C+226%2C+245%2C+280%7D lets just say b is good and call it a day :/

OpenStudy (amistre64):

235.5 -145 ------- 90.5 i cant average for nothing at the moment

OpenStudy (anonymous):



OpenStudy (anonymous):

Idk how to accurately get answer but it is B